# Fan distributions via Tverberg partitions and Gale duality

### v2

@@ Abstract.
Equipartition theory, beginning with the classical ham sandwich theorem, seeks the fair division of finite point sets in $\mathbb{R}^{d}$ by the full-dimensional regions determined by a prescribed geometric dissection of $\mathbb{R}^{d}$. Here we examine equidistributions of finite point sets in $\mathbb{R}^{d}$ by prescribed low dimensional subsets. Our main result states that if $r\geq 3$ is a prime power, then for any $m$-coloring of a sufficiently small point set $X$ in $\mathbb{R}^{d}$, there exists an $r$-fan in $\mathbb{R}^{d}$ – that is, the union of $r$ “half-flats” of codimension $r-2$ centered about a common $(r-1)$-codimensional affine subspace – which captures all the points of $X$ in such a way that each half-flat contains at most an $r$-th of the points from each color class. The number of points in $\mathbb{R}^{d}$ we require for this is essentially tight when $m\geq 2$. Additionally, we extend our equidistribution results to ”piercing” distributions in a similar fashion to Dolnikov’s hyperplane transversal generalization of the ham sandwich theorem. By analogy with recent work of Frick et al., our results are obtained by applying Gale duality to linear cases of topological Tverberg-type theorems. Finally, we extend our distribution results to multiple $r$-fans after establishing a multiple intersection version of a topological Tverberg-type theorem due to Sarkaria.
@@ 1.1. From Equipartitions to Equidistributions of Colored Point Sets
In recent years, equipartition problems have come to occupy a central place in geometric and topological combinatorics. Given a fixed integer $r\geq 2$ and a finite collection of finite point sets (or finite continuous measures) in $\mathbb{R}^{d}$, one seeks a prescribed geometric division of $\mathbb{R}^{d}$ into $r$ interior disjoint and typically convex regions, each of whose interiors contains no more than $1/r$ of the total number of points from each point set (respectively, of each total measure). The prototypical such result is the classical ham sandwich theorem [5, 31], which asserts that any $d$ finite points sets (or finite continuous measures) in $\mathbb{R}^{d}$ can be equipartitioned by a hyperplane, that is, by the two open half-spaces determined by the hyperplane. We refer the reader to the excellent recent survey [26] for more on equipartition theory.
@@
Instead of equipartitions of finite point sets in $\mathbb{R}^{d}$ by full dimensional regions, here we seek their equidistribution by small-dimensional subsets of $\mathbb{R}^{d}$ arising from families of hyperplane with prescribed intersections. Specifically, suppose that $1\leq k\leq\min\{d-1,r-1\}$ and that $H_{1},\ldots,H_{r}$ are hyperplanes in $\mathbb{R}^{d}$ such that (1) the $r$-fold intersection of the hyperplanes is a flat of codimension $k+1$ and (2) any $k+1$ of the $r$ hyperplanes are independent (i.e., their normal vectors are linearly independent). Taking indices modulo $r$, for each $j\in[r]$ let $A_{j}$ be the affine space in $\mathbb{R}^{d}$ of codimension $k$ given by the $k$-fold intersection $A_{j}=\cap_{i=j+1}^{j+k}H_{i}$ of the hyperplanes and let $B_{j}=H_{j}^{+}\cap A_{j}$ be one of the (closed) half-flats of $A_{j}$ determined by a half-space $H_{j}^{+}$ of $H_{j}$. We call the union $F_{r}=\cup_{j=1}^{r}B_{j}$ of these half-flats an $r$-fan of codimension $k$. The boundary of each half-flat is the common intersection $C=\cap_{i=1}^{r}H_{i}$ of the $r$ hyperplanes, which we call the center of the fan. In particular, an $r$-fan of codimension one is the union of $r$ half-hyperplanes centered about an affine space of codimension two. Equipartitions by the regions determined by such $r$-fans have been previously considered (see, e.g., [2, 3, 30]). We say that an $r$-fan of codimension one is regular if the dihedral angle between each of its successive half-hyperplanes is equal to $2\pi/r$.
@@
Now suppose that $X$ is a finite set in $\mathbb{R}^{d}$. We say that an $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{R}^{d}$ distributes X if it contains $X$. If $X=X_{1}\sqcup\cdots\sqcup X_{m}$ is a partition of $X$ into $m$ non-empty subsets, which we call an $m$-coloring of $X$, we say that an $r$-fan equidistributes the coloring if it distributes $X$ and if the interior $\operatorname{\mathrm{Int}}(B_{j})=B_{j}\setminus C$ of each half-flat contains at most $1/r$ of the points from each $X_{i}$.
@@
We have the following equidistribution result for $r$-fans of codimension $r-2$. Recall that a set of points in $\mathbb{R}^{d}$ is affinely spanning if it is not contained in a hyperplane.
@@
Let $d,m\geq 1$ be integers, let $r\geq 3$ be a prime power, and let $n\geq(r-1)(d+m+1)+1$. Then any $m$-coloring of a set of $n$ affinely spanning points in $\mathbb{R}^{n-d-1}$ can be equidistributed by an $r$-fan of codimension $r-2$.
Our proof of Theorem 1.1 shows that the interior of each half-flat of the equidistributing $r$-fan must contain at least one point from $X$, and, as we show in Section 5.1, for typical point sets the interiors of the half-flats must share at least $(r-1)(d+1)+1$ points of $X$. Similar remarks apply to Theorems 1.3, 1.4, and 1.6 below. As an example, letting $r=4$ and $d=m=1$ in Theorem 1.1 shows that for a typical set $X$ of $10$ points in $\mathbb{R}^{8}$ there is a $6$-dimensional $4$-fan which catches all $10$ points of $X$ so that either (a) $2$ points of $X$ lie in the fan’s $5$-dimensional center while the remaining $8$ points are equally dividing among the fan’s $4$ open half-flats, or (b) $3$ points lie in the center while the remaining $7$ are divided among its open half-flats so that each receives at most $2$ points.
Proposition 5.3 below will show that Theorem 1.1 is nearly sharp with respect to dimension for all fixed $r\geq 3$ and $m\geq 2$. Namely, for each integer $d\geq r-1$, let $n(r,m,d)$ be the maximum integer $n$ such that any $m$-coloring of any affinely spanning set of $n$ points in $\mathbb{R}^{d}$ can be equidistributed by an $r$-fan of codimension $r-2$. Once $d$ is sufficiently large, the lower bound on $n(r,m,d)$ given by Theorem 1.1 exceeds the exact value by at most two:
@@
Let $r\geq 3$ be a prime power, let $m\geq 2$ be an integer, let $c=(r-1)(m+1)$, and let $d$ be an integer of the form $d=(r-2)s+t+c$ for some $s\geq 1$ and $t\in\{0,\ldots,r-2\}$. Then
$d+s\leq n(r,m,d)\leq d+s+2.$
When $r=3$, the $r$-fans of Theorem 1.1 have codimension one but need not be regular. If one considers point sets in $\mathbb{C}^{d}$, however, then for any prime power $r$ one can guarantee equidistributions by a complex regular $r$-fan in $\mathbb{C}^{d}$, that is a regular $r$-fan in $\mathbb{R}^{2d}$ of codimension one whose center is a complex hyperplane. For this, we require that the points complex affinely span $\mathbb{C}^{d}$, or in other words that they do not all lie on a complex hyperplane.
@@
As a regular $2$-fan in $\mathbb{R}^{d}$ is a hyperplane, the $m=2$ case of Theorem  1.3 is closely aligned with the ham sandwich theorem. Namely, let $d=s+m+1+t$, where $s\geq 1$ and $t\in\{0,1\}$. Then for any $m$-coloring $X=X_{1}\sqcup\cdots\sqcup X_{m}$ of a set $X$ of $n:=d+s+1$ complex affinely spanning points in $\mathbb{C}^{d}$, there is a real hyperplane $H$ in $\mathbb{C}^{d}$ containing $X$ and a complex hyperplane $H_{\mathbb{C}}$ contained in $H$ such that $H_{\mathbb{C}}$, when viewed as a real hyperplane in $H$, equipartitions each of the $X_{i}$. The number of points in $\mathbb{C}^{d}$ required for this is exceeds the optimum number by at most $2$ for any $m\geq 1$ (see Remark 5.4 below).
@@ 1.2. Piercing Distributions
Recall that a family $\mathcal{F}$ of subsets of a finite set $X$ in $\mathbb{R}^{d}$ is said to be intersecting if any two members of $\mathcal{F}$ have non-empty intersection. The discrete case of a well-known theorem of Dolnikov [9] states that if $\mathcal{F}_{1},\ldots,\mathcal{F}_{d}$ are intersecting families of subsets of a finite set in $\mathbb{R}^{d}$, then there exists a hyperplane in $\mathbb{R}^{d}$ which intersects the convex hull of every member of each family. This result easily implies the ham sandwich theorem (see, e.g., [9, 12]). For prime powers $r\geq 3$, we give distribution results which are similar in spirit to Dolnikov’s theorem under the weaker intersection criterion that no family contains $r$ pairwise disjoint sets. Instead of a hyperplane which pierces the convex hull of every member of each family, we are guaranteed a distributing $r$-fan such every member of each family is itself pierced by two of the fan’s half-flats.
In what follows, it will be convenient to reformulate our intersection condition in terms of hypergraph colorings. Given a finite family $\mathcal{F}$ of non-empty subsets of a set $X$, its $r$-uniform Kneser hypergraph $\mathrm{KG}^{r}(\mathcal{F})$ is the $r$-uniform hypergraph whose vertices are the elements of $\mathcal{F}$ and whose hyperedges $\{F_{1},\ldots,F_{r}\}$ consist of $r$ pairwise-disjoint elements from $\mathcal{F}$. The chromatic number $\chi(\mathrm{KG}^{r}(\mathcal{F}))$ of $\mathrm{KG}^{r}(\mathcal{F})$ is the minimum number of colors needed to color all the vertices of $\mathrm{KG}^{r}(\mathcal{F})$ so that no hyperedge is monochromatic. Thus to say that $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$ means that there are $m$ non-empty subfamilies $\mathcal{F}_{1},\ldots,\mathcal{F}_{m}$ of $\mathcal{F}$ such that $\mathcal{F}=\cup_{j=1}^{r}\mathcal{F}_{i}$ and no $\mathcal{F}_{i}$ contains $r$ pairwise disjoint sets.
@@
    Suppose that $r\geq 3$ and that $X$ is a set of $n\geq(r-1)(d+1)+1$ affinely spanning points in $\mathbb{R}^{n-d-1}$. If $\mathcal{F}$ is a family of non-empty subsets of $X$ with $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$, then there is an $r$-fan of codimension $r-2$ which distributes $X$ so that every $F\in\mathcal{F}$ is intersected by at least two closed half-flats of $F_{r}$.
@@
    Suppose that $r\geq 2$ and that $X$ is a set of $n\geq(r-1)(2d+m+1)+1$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$. If $\mathcal{F}$ is a family of non-empty subsets of $X$ with $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$, then there is a complex regular $r$-fan $F_{r}$ in $\mathbb{C}^{n-d-1}$ which distributes $X$ so that every $F\in\mathcal{F}$ is intersected by at least two closed half-hyperplanes of $F_{r}$.
As an example of Theorem 1.4(a), letting $r=4$ and $d=1$ shows that for any affinely spanning set $X$ of $10$ points in $\mathbb{R}^{8}$, there is a $6$-dimensional $4$-fan in $\mathbb{R}^{8}$ which contains $X$ so that none of the three element subsets of $X$ is contained in the interior of any of the fan’s half-flats. If $X$ is typical, then the interiors of the fan’s half-flats must share at least $7$ of the $10$ points.
The $r=2$ case of Theorem 1.4(b) gives an analogue of Dolnikov’s theorem for complex instead of real hyperplanes: If $d=s+m+1+t$ with $s\geq 1$ and $t\in\{0,1\}$, let $X$ be an affinely spanning set of $n:=d+s+1$ points in $\mathbb{C}^{d}$ and let $\mathcal{F}_{1},\ldots,\mathcal{F}_{m}$ be non-empty intersecting families of subsets of $X$. Then exists a real hyperplane $H$ in $\mathbb{C}^{d}$ which contains $X$ and a complex hyperplane $H_{\mathbb{C}}$ contained in $H$ such that $H_{\mathbb{C}}$ intersects the convex hull of every member of each $\mathcal{F}_{i}$.

### v3

@@ Abstract.
Equipartition theory, beginning with the classical ham sandwich theorem, seeks the fair division of finite point sets in $\mathbb{R}^{d}$ by the full-dimensional regions determined by a prescribed geometric dissection of $\mathbb{R}^{d}$. Here we examine equidistributions of finite point sets in $\mathbb{R}^{d}$ by prescribed low dimensional subsets of $\mathbb{R}^{d}$. Our main result states that if $r\geq 3$ is a prime power, then for any $m$-coloring of a sufficiently small point set $X$ in $\mathbb{R}^{d}$, there exists an $r$-fan in $\mathbb{R}^{d}$ – that is, the union of $r$ “half-flats” of codimension r2r-2 centered about a common (r1)(r-1)-codimensional affine subspace – which captures all the points of $X$ in such a way that each half-flat contains at most an $r$-th of the points from each color class. The number of points in $\mathbb{R}^{d}$ we require for this is essentially sharp when $m\geq 2$ and in particular is asymptotically tight. Additionally, we extend our equidistribution results to “piercing” distributions in a similar fashion to Dolnikov’s hyperplane transversal generalization of the ham sandwich theorem. By analogy with recent work of Frick et al., our results are obtained by applying Gale duality to linear cases of topological Tverberg-type theorems. Finally, we extend our distribution results to multiple $r$-fans after establishing a multiple intersection version of a topological Tverberg-type theorem due to Sarkaria.
@@ 1.1. From Equipartitions to Equidistributions of Colored Point Sets
In recent years, equipartition problems have come to occupy a central place in geometric and topological combinatorics. Given a fixed integer $r\geq 2$ and a finite collection of finite point sets (or finite continuous measures) in $\mathbb{R}^{d}$, one seeks a prescribed geometric division of $\mathbb{R}^{d}$ into $r$ interior disjoint and typically convex regions, each of whose interiors contains no more than $1/r$ of the total number of points from each point set (respectively, of each total measure). The prototypical such result is the classical ham sandwich theorem [6, 34], which asserts that any $d$ finite points sets (or finite continuous measures) in $\mathbb{R}^{d}$ can be equipartitioned by an affine hyperplane, that is, by the two half-spaces determined by the hyperplane. We refer the reader to the excellent recent survey [27] for more on equipartition theory.
@@
Instead of equipartitions of finite point sets in $\mathbb{R}^{d}$ by full dimensional regions, here we seek their equidistribution by low dimensional subsets. Specifically, suppose that $A$ is a $k$-flat in $\mathbb{R}^{d}$, where $1\leq k\leq d-1$. Given an affine hyperplane $H$ in $\mathbb{R}^{d}$ which intersects but does not contain $A$, we call the intersection $B=A\cap H^{+}$ of $A$ with a (closed) half-space $H^{+}$ of $H$ a (closed) half $k$-flat. Thus the boundary $\partial B=A\cap H$ of $B$ is a (k1)(k-1)-flat. For $r\geq 2$, we say that the union $F_{r}=\cup_{j=1}^{r}B_{j}$ of $r$ half $k$-flats $B_{1},\ldots,B_{r}$ is a $k$-dimensional $r$-fan provided that there is a single (k1)(k-1)-flat $C$, called the center of the fan, which is the boundary of each $B_{j}$. A (dk)(d-k)-dimensional $r$-fan will be said to have codimension $k$. In particular, an $r$-fan of codimension one is the union of $r$ half-hyperplanes centered about an affine space of codimension two. Equipartitions by the regions determined by such $r$-fans have been previously considered (see, e.g., [3, 4, 32]).
@@
Now suppose that $X$ is a finite set in $\mathbb{R}^{d}$. We say that an $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{R}^{d}$ distributes X if it contains $X$. If $X=X_{1}\sqcup\cdots\sqcup X_{m}$ is a partition of $X$ into $m$ non-empty subsets, which we call an $m$-coloring of $X$, we say that an $r$-fan equidistributes this coloring if it distributes $X$ so that the interior $\operatorname{\mathrm{Int}}(B_{j})=B_{j}\setminus C$ of each half-flat contains at most $1/r$ of the points from each $X_{i}$.
We present equidistribution results for two types of $r$-fans, the first of which is given by certain hyperplane arrangements. Namely, suppose that $H_{1},\ldots,H_{r}$ are oriented affine hyperplanes in $\mathbb{R}^{d}$ such that (1) any r1r-1 of the $r$ hyperplanes are independent (i.e., their normal vectors are linearly independent) and (2) the $r$ hyperplanes are dependent and the $r$-fold intersection of the hyperplanes is non-empty (and so is a flat of codimension r1r-1). Taking indices cyclically, for each $j\in[r]$ we let $A_{j}=\cap_{i\neq j-1,j}H_{i}$ be the (r2)(r-2)-codimensional flat given by intersecting all of the hyperplanes except $H_{j-1}$ and $H_{j}$, and we let $B_{j}=H_{j}^{+}\cap A_{j}$ be the half-flat given by intersecting $A_{j}$ with the half-space $H_{j}^{+}$ of $H_{j}$. The union $F_{r}=\cup_{j=1}^{r}B_{j}$ is then an $r$-fan of codimension r2r-2 whose (r1)(r-1)-codimensional center is the common intersection of the $H_{i}$. We shall say that such an $r$-fan is conical (See Figure 1 below for a picture when $r=3$). For instance, a conical $r$-fan in $\mathbb{R}^{r-1}$ is a union of $r$ rays emanating from a common point. Our second class of fans are regular $r$-fans. These are fans of codimension one for which the dihedral angle between any two successive half-hyperplanes is $2\pi/r$.
@@
We have the following equidistribution result for conical $r$-fans. Recall that a set of points in $\mathbb{R}^{d}$ is affinely spanning if it is not contained in an affine hyperplane.
@@
Let $d,m\geq 1$ be integers, let $r\geq 3$ be a prime power, and let $n\geq(r-1)(d+m+1)+1$. Then any $m$-coloring of a set of $n$ affinely spanning points in $\mathbb{R}^{n-d-1}$ can be equidistributed by a conical $r$-fan.
Our proof of Theorem 1.1 shows that the interior of each half-flat of the conical $r$-fan must contain at least one point from $X$, and, as we show in Section 8.1, for “typical” point sets the union of the interiors of the half-flats contains at least (r1)(d+1)+1(r-1)(d+1)+1 points of $X$. Similar remarks apply to Theorems 1.3, 1.4, and 1.6 below. As an example, letting $r=3$ and $d=m=1$ in Theorem 1.1 shows that for a typical set $X$ of $7$ points in $\mathbb{R}^{5}$ there is a ($4$-dimensional) conical $3$-fan in $\mathbb{R}^{5}$ which catches all $7$ points of $X$ so that either (a) $1$ point of $X$ lies in the fan’s $3$-dimensional center while the remaining $6$ points are equally divided among the fan’s $3$ open half-flats, or (b) $2$ points lie in the center while the remaining $5$ are divided so that $2$ of the open half-flats receive $2$ points each and the remaining open half flat receives a single point. Figure 1 gives a picture of case (a) when viewed in the quotient $\mathbb{R}^{5}/C\cong\mathbb{R}^{2}$ of $\mathbb{R}^{5}$ by the fan’s center $C$.
Figure 1. Equidistribution of $7$ points by a conical $3$-fan
Proposition 8.4 below will show in particular that Theorem 1.1 is nearly sharp with respect to dimension for all fixed $r\geq 3$ and $m\geq 2$ and so is asymptotically tight. Namely, for each integer $d\geq r-1$ let $n(r,m,d)$ be the maximum integer $n$ such that any $m$-coloring of any affinely spanning set of $n$ points in $\mathbb{R}^{d}$ can be equidistributed by a conical $r$-fan. Once $d$ is sufficiently large, the lower bound on $n(r,m,d)$ given by Theorem 1.1 exceeds the exact value by at most four:
@@
Let $r\geq 3$ be a prime power and let $m\geq 2$ be an integer. Suppose that $d$ is an integer of the form d=(r2)s+t+cd=(r-2)s+t+c, where $s\geq 1$, $t\in\{0,\ldots,r-3\}$, and c=(r1)(m+1)c=(r-1)(m+1). Then
$d+s+1\leq n(r,m,d)\leq d+s+5.$
When $r=3$, the $r$-fans of Theorem 1.1 have codimension one but need not be regular. If one considers point sets in $\mathbb{C}^{d}$, however, then for any prime power $r$ one can guarantee equidistributions by a complex regular $r$-fan in $\mathbb{C}^{d}$, that is a regular $r$-fan in $\mathbb{R}^{2d}$ whose center is a complex affine hyperplane. For this we require that the points complex affinely span $\mathbb{C}^{d}$, or in other words that they do not all lie on a complex affine hyperplane. See [28] for a related complex analogue of the well-known central transversal theorem.
@@
As a regular $2$-fan in $\mathbb{R}^{d}$ is an affine hyperplane, the $m=2$ case of Theorem  1.3 is closely aligned with the ham sandwich theorem. Namely, let $d=s+m+1$, where $s\geq 1$. Then for any $m$-coloring $X=X_{1}\sqcup\cdots\sqcup X_{m}$ of a set $X$ of $d+s+1$ complex affinely spanning points in $\mathbb{C}^{d}$, there is a real affine hyperplane $H$ in $\mathbb{C}^{d}$ containing $X$ and a complex affine hyperplane $H_{\mathbb{C}}$ contained in $H$ such that $H_{\mathbb{C}}$ – when viewed as a real affine hyperplane in $H$ – equipartitions each of the $X_{i}$. The number of points in $\mathbb{C}^{d}$ required for this is almost tight for all $m\geq 1$ (see Remark 8.5).
@@ 1.2. Piercing Distributions
Recall that a family $\mathcal{F}$ of non-empty subsets of a finite set $X$ in $\mathbb{R}^{d}$ is said to be intersecting if any two members of $\mathcal{F}$ have non-empty intersection. The discrete case of a well-known theorem of Dolnikov [10] states that if $\mathcal{F}_{1},\ldots,\mathcal{F}_{d}$ are intersecting families of non-empty subsets of a finite set in $\mathbb{R}^{d}$, then there exists an affine hyperplane in $\mathbb{R}^{d}$ which intersects the convex hull of every member of each family. This result easily implies the ham sandwich theorem (see, e.g., [10, 13]). For prime powers $r\geq 3$, we give distribution results which are similar in spirit to Dolnikov’s theorem under the weaker intersection criterion that no family contains $r$ pairwise disjoint sets. Instead of an affine hyperplane which pierces the convex hull of every member of each family, we are guaranteed a distributing $r$-fan such that every member of each family is itself pierced by two of the fan’s half-flats.
In what follows, it will be convenient to reformulate our intersection condition in terms of hypergraph colorings. Given a finite family $\mathcal{F}$ of non-empty subsets of a set $X$, its $r$-uniform Kneser hypergraph $\mathrm{KG}^{r}(\mathcal{F})$ is the $r$-uniform hypergraph whose vertices are the elements of $\mathcal{F}$ and whose hyperedges $\{F_{1},\ldots,F_{r}\}$ consist of $r$ pairwise-disjoint elements from $\mathcal{F}$. The chromatic number $\chi(\mathrm{KG}^{r}(\mathcal{F}))$ of $\mathrm{KG}^{r}(\mathcal{F})$ is the minimum number of colors needed to color all the vertices of $\mathrm{KG}^{r}(\mathcal{F})$ so that no hyperedge is monochromatic. Thus to say that $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$ means that there are $m$ non-empty subfamilies $\mathcal{F}_{1},\ldots,\mathcal{F}_{m}$ of $\mathcal{F}$ such that $\mathcal{F}=\cup_{j=1}^{m}\mathcal{F}_{i}$ and no $\mathcal{F}_{i}$ contains $r$ pairwise disjoint sets.
@@
    Suppose that $r\geq 3$ and that $X$ is a set of $n\geq(r-1)(d+1)+1$ affinely spanning points in $\mathbb{R}^{n-d-1}$. If $\mathcal{F}$ is a family of non-empty subsets of $X$ with $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$, then there is a conical $r$-fan which distributes $X$ so that every $A\in\mathcal{F}$ is intersected by at least two closed half-flats of $F_{r}$.
@@
    Suppose that $r\geq 2$ and that $X$ is a set of $n\geq(r-1)(2d+m+1)+1$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$. If $\mathcal{F}$ is a family of non-empty subsets of $X$ with $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$, then there is a complex regular $r$-fan $F_{r}$ in $\mathbb{C}^{n-d-1}$ which distributes $X$ so that every $A\in\mathcal{F}$ is intersected by at least two closed half-hyperplanes of $F_{r}$.
The $r=2$ case of Theorem 1.4(b) gives an analogue of Dolnikov’s theorem for complex instead of real hyperplanes: If $d=s+m+1$ with $s\geq 1$, let $X$ be an affinely spanning set of $d+s+1$ points in $\mathbb{C}^{d}$ and let $\mathcal{F}_{1},\ldots,\mathcal{F}_{m}$ be non-empty intersecting families of subsets of $X$. Then exists a real affine hyperplane $H$ in $\mathbb{C}^{d}$ which contains $X$ and a complex affine hyperplane $H_{\mathbb{C}}$ contained in $H$ such that $H_{\mathbb{C}}$ intersects the convex hull of every member of each $\mathcal{F}_{i}$.